OpenStudy (anonymous):

please need answer quick x^2-2x-10=0

OpenStudy (anonymous):

\[(x-5)(x+2)\]

OpenStudy (anonymous):

(x-5)(x+2) = x^2 - 3x - 10 so you will have to use the quadratic equation: \[x = 1 \pm \sqrt{11}\]